OpenStudy (helpjebal):

Solve by Factoring: 4z^2 - 4z +1 = 0

OpenStudy (helpjebal):

I'm having a little trouble getting past the first step. So far I have: 4z^2 - 4z = -1 but maybe it's supposed to start off with 4(z^2 - z) +1 = 0?

OpenStudy (solomonzelman):

you have to factor the equation... (Hint: you have a perfect square polynomial in the form of a²-2ab+b²=(a-b)² )

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 4z^2-4z+1\quad\Longrightarrow\quad (2z)^2-2(2z)(1)+(1)^2 }\) \(\tiny{\\[1.5em]}\) \(\color{#000000 }{ \displaystyle (2z)^2-2(2z)(1)+(1)^2\quad\Longrightarrow\quad (2z-1)^2}\)\(\tiny{\\[1.3em]}\) \(\color{#000000 }{ \displaystyle (a)^2-2(a)(b)+(b)^2\quad\Longrightarrow\quad (a-b)^2}\)

OpenStudy (solomonzelman):

So, \(\color{#000000 }{ \displaystyle 4z^2-4z+1=(2z-1)^2 }\)

OpenStudy (solomonzelman):

Now, you have, \(\color{#000000 }{ \displaystyle (2z-1)^2=0 }\) Solve for \(\color{black}{z}\)...

OpenStudy (helpjebal):

2z = 1 z = 1/2

OpenStudy (helpjebal):

thanks!